OpenStudy (anonymous):

Find the product. (n^3)^2· (n^5)^4

OpenStudy (anonymous):

n=?

OpenStudy (anonymous):

n^(3x2). n^(5x4) = n^6. n^20 = n^(6+20) = n^ 26